OpenStudy (anonymous):

Please help! The polar coordinates (8, 7π/4) represent the Cartesian point a. (4, 4) b. (4√2, 4√2) c. (-4√2, 4√2) d. (4√2, -4√2)

OpenStudy (anonymous):

\[ x= \rho \cos(\theta)\\ y= \rho \sin(\theta) \]

OpenStudy (anonymous):

Use \[ \rho= 8\\ \theta= 7\pi/4 \]

OpenStudy (anonymous):

Notice that \[7\pi/4 = 2 \pi - \pi/4 \]

OpenStudy (anonymous):

Are you there?

OpenStudy (anonymous):

i've got x= 8cos(7pi/4) and y=8sin(7pi/4)

OpenStudy (anonymous):

How much is the cos and the sin of your angle and you will be done.

OpenStudy (anonymous):

\[ \cos(\pi/4) = \sin(\pi/4)= \frac { \sqrt 2}2 \]

OpenStudy (anonymous):

It is easy to find your result now.

OpenStudy (anonymous):

\[ \cos( 2 \pi -\theta) = \cos(-\theta) =\cos(\theta) \]

OpenStudy (anonymous):

\[ \sin( 2 \pi -\theta) = \sin(-\theta) =-\sin(\theta) \]

OpenStudy (anonymous):

Can you finish it now?